What is the probability of rolling at least two sixes when a fair six sided die is rolled 5 times?
It is 0.1962
What are the possible outcomes for ABCD?
The answer will depend on what ABCD is (or are). And since you have not bothered to provide that information, I cannot provide a sensible answer.
The king comes from a family of three children What is the probability that he has a brother?
The probability of a boy (male) is equal to the probability of a girl (female) which equals 1/2. The king is a male. So, we need the probability of a male and a male which is 1/2 * 1/2 = 1/4.
When is theoretical less than the experimental?
theoretical probability can be smaller than expiremental like this
say you toss a coin 4 times
it lands on head 3 times
that means the theoretical 1/2 is now smaller than the expiremental 3/4
What is the value of n when 6 over n times 5 over n-1 equals 1 over 3 showing all work?
If you mean: 6/n times 5/n-1 = 1/3
Then: 30/n2-n = 1/3
Multiplying both sides by n2-n: 30 = n2-n/3
Multiplying both sides by 3: 90 = n2-n
Subtracting 90 from both sides: 0 = n2-n-90 or n2-n-90 = 0
Solving the above quadratic equation: n = -9 or n =10
If n is of a material value its more likely to be 10
Note that n2 means n squared
What does distribution of coal mean?
It could mean the geological spread of coal seams in the ground. Or it could mean the process of getting the coal from the mines to the consumers.
What is your probability of getting hired as a neonatologist in the future?
The probability is 0 since I have no interest, not aptitude, in that profession.
How much sand is needed for a 16 X 32 pool?
Simply ask the company who would be delivering your sand for the sized pool. They usually know the average cubic tons per size pool. Should be 4 tons.
There is only one girl out of 12 students so the probability that the girl is selected is 1/12.
How to identify the the shape of probability distribution?
You cannot. There are hundreds of different distributions. The shapes of the distributions depend on their parameters so that the same distribution can be symmetric when the parameters have some specific value, but is highly skewed - in either direction - for other values.
Using 5 and 6 what can you determine about probability and actual answer?
Is this another quiz question? You are missing half the question.
What is the probability of rolling a number greater than or equal to 3?
The answer depends on what you are rolling: a number cube or some other shape? For a die, the answer is 2/3.
How do you work out probability questions?
You divide (the number of ways the try can succeed) by (the total number of ways it can come out).
How many ways can you arrange 5 items if two have to be next to each other?
There are 48 possibilities where 2 of 5 items must be adjacent.
If you have 5 items, and they can be arranged in any order, there are 120
5 x 4 x 3 x 2 ways to arrange them, for example ABCDE, ABCED, and so forth.
However, if any two need to be placed next to each other, the number of variations is reduced to [4 x 3 x 2] x2 (=48), where there are only 4 separate "units" arranged, but the double-unit can appear with either of the pair first.
For example, if A and B must be together, you have 24 possibilities:
(AB)CDE
(AB)CED
(AB)DCE
(AB)DEC
(AB)EDC
(AB)ECD
C(AB)DE
C(AB)ED
D(AB)CE
D(AB)EC
E(AB)CD
E(AB)DC
CD(AB)E
DC(AB)E
CE(AB)D
EC(AB)D
DE(AB)C
ED(AB)C
CDE(AB)
CED(AB)
DEC(AB)
DCE(AB)
EDC(AB)
ECD(AB)
and another 24 where (AB) is replaced by (BA).
What is the probability of 5 winning numbers between one and 31?
If the winning numbers are picked at random, the probability is 1 in 169911.
It is to celebrate the day of your birth. and the reason why peoples 18th birthday is so important is because they are growing into an adult.
What are the odds of having kids with the same birthday but a different year?
In probability theory, the birthday problem, or birthday paradox[1] pertains to the probability that in a set of randomly chosen people some pair of them will have the same birthday. In a group of 10 randomly chosen people, there is an 11.7% chance. In a group of at least 23 randomly chosen people, there is more than 50% probability that some pair of them will both have been born on the same day. For 57 or more people, the probability is more than 99%, and it reaches 100% when the number of people reaches 367 (there are a maximum of 366 possible birthdays). The mathematics behind this problem leads to a well-known cryptographic attack called the birthday attack. See Wikipedia for more: http://en.wikipedia.org/wiki/Birthday_paradox
Undefying end was a word god told me He said he would bring this to an undefying end!
What is the probability of picking any red coloured card from a pack?
it is 13/52 because there are always 4 different card colours in each deck so divide 4 in 52 and your answer is 13
hope you understand my equation
* * * * *
Hope you don't because it is incorrect! The probability of any event is a number in the range [0, 1] so the answer cannot be 13!
A normal deck two coulours, not 4: red and black. It therefore contains 26 red cards and so the probability is 26/52 = 1/2.
False; the "or" is an additive property so the probability of rain or snow muse be greater than or equal to 0.65.